OpenStudy (anonymous):

a model rocket is launched from a roof into a large field. The path of the rocket can be modeled by the equation y=-0.04x^2 +8.3x +4.3 , where x is the horizontal distance, in meters, from the starting point on the roof and y is the height, in meters, of the rocket above the ground. how far horizontally from its starting point will the rocket land? Round your answer to the nearest hundredth meter.

OpenStudy (campbell_st):

well you need to let y = 0 that means the reockat has no height, and solve for x easiest way seems to be the general quadratic formula \[x = \frac{-8.3 \pm \sqrt{8.3^2 - 4 \times -0.04 \times 4.3}}{2 \times -0.04}\] hope this helps.

OpenStudy (anonymous):

so all I had to do was plug a b and c into the quadratic formula??

OpenStudy (campbell_st):

thats correct... the picture looks like  and you will only have 1 answer, the positive value of x